Supose the consumer`s utility function is given by U= 10X1/2Y1/4. Compute both MRSxy and MRSyx. Is there any difference between the two?



1
Expert's answer
2021-08-05T13:06:53-0400

Solution:

U = 10X1/22Y1/4

MRSxy = "\frac{MUx}{MUy}"


MUx = "\frac{\partial U} {\partial X}" = 5X-1/22Y1/4


MUy = "\frac{\partial U} {\partial Y}" = 2.5X1/22Y-3/4


MRSxy = "\frac{5X^{-1/2}2Y^{1/4} } {2.5X^{1/2}2Y^{-3/4}}" = "\frac{5Y } {2.5X}"


MRSxy = "\frac{5Y } {2.5X}"

 

 

MRSyx = "\frac{MUy}{MUx}"


MUy = "\frac{\partial U} {\partial Y}" = 2.5X1/22Y-3/4


MUx = "\frac{\partial U} {\partial X}" = 5X-1/22Y1/4


MRSyx = "\frac{2.5X^{1/2}2Y^{-3/4} } {5X^{-1/2}2Y^{1/4}} = \frac{2.5X } {5Y}"


MRSyx = "\frac{2.5X } {5Y}"

 

Both MRS are different since they are showing different values and hence will produce different results.